# A Novel Energy-Efficient Cross-Layer Design for Scheduling and Routing in 6TiSCH Networks

### v1

@@ III-D Scheduling Information Injection
In the context of TSCH, transmitting packets from higher-layer protocols is not an instantaneous process; rather, it entails a meticulous queuing and scheduling mechanism. As depicted in Fig. 7, incoming packets are first enqueued within the TSCH queue and await their transmissions. Since we inject the proposed time slots in DIO and DAO packets, they will be locked during time $T_{q}=T_{\rm de}-T_{\rm en}$, where $T_{\rm en}$ is the instant at which the packet entered the TSCH queue, and $T_{\rm de}$ is the time at which it left the queue. To avoid locking the proposed time slots for a long time $T_{q}$, we propose to inject them within the packet only before leaving the queue, i.e., at instant $T_{de}$, hence $T_{q}\approx 0$. By carefully orchestrating the data injection in the RPL control packets, we optimize time slot utilization.
@@
III-E Avoiding TSCH Queue Overflow
@@
III-F Adaptive Self-Sufficient Cell Reservation
In a typical 6TiSCH network, all DAO packets are destined towards the DAG root, thus navigating in a multi-hop fashion. In our proposal, essential information about cell reservation, which consists of the available slots and proposed slot lists, is embedded within these DAO packets. When a parent node receives a DAO packet, as it is the principal conduit to the DAG root, it reserves new cells for its new child where the number of cells should not exceed the maximum number of cells to allocate at once. Then, it loads the list of its free slots (instead of the old one that belongs to the child) and the number of slots to be selected (slots to be allocated to its new child) into the DAO packet. This packet is then seamlessly forwarded, with each intermediate node towards the DAG root executing the same procedure. As shown in Fig. 8, upon reception of a DAO packet from node $3$, node $i$ undertakes cell selection for reservation and embeds the list of chosen slots into the TSCH ACK packet. It then forwards the enhanced packet to its parent. To reduce latency at every hop, we propose that the receiver of a DAO packet selects the nearest time slot to the actual one used for the reception. One way to do it is to rely on the concept of the Low-Latency SF (LLSF) protocol [31] where the slots are selected according to the gap to the left of the actual slot. However, we propose here to select the nearest slot to the actual one in any direction. Accordingly, if the node is joining the DODAG, the actual slot will be in the list of proposed slots, thus maximizing its chance of selection. If the node has already joined the DODAG, then the actual slot is the slot that is already allocated to its parent. Consequently, both energy and latency tied to time slot reservations can be reduced.
@@ III-F Adaptive Self-Sufficient Cell Reservation
TABLE I: Simulation parameters
Parameters Value
Number of nodes
50, 100
Number of simulation runs
100
Simulation duration (min)
30, 270
[…]
1
Service or application packet period (sec)
5, 15, 45, 60
Slotframe length (in number of slots)
100
Time slot duration (sec)
0.01
Secure Joining required
True
TSCH maximum payload length (bytes)
120
TSCH queue size (number of packets)
10
DIO packet size (bytes)
MSF
PB
76
120
DAO packet size (bytes)
20
75
IV Performance evaluation
IV-A Simulation configuration
To comprehensively assess the efficiency and resilience of our proposed scheme, we employed the 6TiSCH simulator [32]. For the sake of thorough analysis, generated plots include outlier results, thus reflecting the full range of observed results. Our simulations spanned $30$ minutes and $270$ minutes ($4.5$ hours) and were repeated 100 times across two distinct network configurations:
@@ IV-A Simulation configuration
[…]
TABLE II: Parameters of the proposed solution
Parameter Value
Number of proposed slots per DIO
7
DIO cells duration (in number of slotframes)
10
Number of permanent DIO slots
1
Slot selection ratio within DIO slots
3
Maximum number of cells to reserve at once
5
Initial phase duration (min)
45
TSCH queue empty places threshold for cell reservation
2
Cycle interval before next cell request
2
Number of cells reserved upon reservation process
1
Figure 9: Joining time vs. nbr. of proposed slots per DIO.
Figure 10: Consumed current vs. nbr. of proposed slots per DIO.
To set the parameter “Number of proposed slots per DIO” in Table II to the best value, we conducted several experiments with different values. Specifically, in Figs. 9-10, we evaluated the impact of different numbers of proposed time slots per DIO on the joining time and current consumption (equivalent to energy consumption). These experiments have been conducted over an observation time of $30$ minutes, with an application period of $5$ seconds, and for different network sizes, i.e., $N_{1}=50$ and $N_{2}=100$ nodes. As shown in Fig. 9, for any number of nodes, the joining time reduces when the number of proposed slots is higher. This is expected since more proposed slots mean that a novel node can rapidly be configured and join the DODAG. However, since the number of proposed slots cannot be very high, as it may increase the size of packets, our preference is for the value $7$ since it provides similar results to the system with the value $13$, in particular for $N_{2}$, i.e., for a large network. A similar explanation can be provided for Fig. 10 where the value $7$ achieves low current consumption for the large network. Indeed, a very high value, such as $13$, can result in inefficient idle listening, thus increasing power consumption. Consequently, we recommend the number of proposed slots per DIO to be $7$ as the best value since it is the lowest one that achieves good performances in terms of joining time and consumed power.
@@
Figure 11: Joining time vs. length of service packet period (in sec) (first 30 minutes).
IV-B Simulation Results
The initial phase of the network’s life is known as network formation. This phase is pivotal for establishing the network’s parameters and topology. To demonstrate the effectiveness of our solution in accelerating the network formation timeline and minimizing power consumption [30], we conducted simulations in Figs. 11-12 under the same conditions as the previous figures, while we varied the value of the parameter “Service packet period”. Also, we compared the results of our proposal, denoted PB, with those of the benchmark MSF.
@@ IV-B Simulation Results
Figure 12: Consumed current vs. length of service packet period (in sec) (first 30 minutes).
@@
Figure 13: Total nbr. of sent/received packets (different service packet periods, 4.5 hours).
As shown in Fig 11, our proposed solution PB achieves lower joining time for any $N$ and service packet period duration. For instance, PB achieves an average joining time of 13 minutes for a service packet period equal to $15$ and $N=100$, while MSF needs approximately 14 minutes in the same conditions. Indeed, a node within the PB framework uses proposed slots in the received DIO to define the cells during which it sends the DAO, unlike MSF which relies on sending a 6P request to its preferred parent and then waiting for a response for cell reservation. In PB, the DAO packet serves a dual purpose: Communicating with the RPL to establish the DODAG structure and participating in SF. As the service packet period increases, we notice a degradation in the joining time. This is expected since a node with a low (resp. high) service packet period will transmit EB packets more (resp. less) often, which increases (resp. reduces) the likelihood that other nodes synchronize and join the DODAG. In 6TiSCH networks, the process of network formation often expends a significant amount of energy, even before useful data transmission. In Fig. 12, we present the consumed current related to the scenarios of Fig.11. The PB scheme performs better than the MSF one, for any $N$ and service packet period duration. The performance gap increases as $N$ becomes higher. This result proves the robustness of our scheme and its superiority when operating in dense networks. Finally, we notice that the service packet period has a slight impact on the total consumed current. After validating the improvements of our solution on network formation, we assess the performances of the schemes for long-term operation. We evaluate the performances in terms of the number of packets transmitted/received, the end-to-end (E2E) latency, network jitter, and consumed charge. For the next results, experiments have been conducted for 4.5 hours. Fig. 13 illustrates the number of transmitted (Tx) and received (Rx) packets for different service packet periods and $N$, where PB and MSF are compared. For any $N$ and service packet period, our solution outperforms the MSF benchmark. In other words, the network that uses our approach realizes more frequent transmissions, and nodes within it successfully receive more packets. When the service packet period increases, i.e., a longer service is running, the number of exchanged packets drops. such a result is expected since during the observation time of 4.5 hours, a longer service time impacts the frequency of its execution over the observation period. Moreover, when $N=100$, the number of exchanged packets almost doubles compared to the case with $N=50$. Indeed, the availability of more nodes joining the DODAG makes it possible to transmit/receive more packets, within the limits of available channels and time slots in 6TiSCH.
@@
Figure 14: E2E latency vs. service packet period (in sec) (4.5 hours).
@@
Figure 15: Median and avg. jitter for transmitted packets during 4.5 hours (different service packet periods).
Fig. 14 shows the network’s E2E latency for different service packet periods and $N$. End-to-end latency is defined as the duration measured from the ASN at which a source node transmits a service/application packet, to the ASN at which this packet is successfully received and acknowledged by the destination. As it can be seen, the PB scheme achieves lower E2E latency in any network condition than the MSF benchmark. This result underscores the efficiency of our method. Even when the service packet period increases, the E2E latency follows in a logarithmic trend with values stabilizing at high values. For instance, for $N=100$, the average E2E latency is almost the same, about 2.1 sec, for service packet periods $45$ and $60$.
In Fig. 15, we plot the median and average jitter performance for the transmitted packets during the observation time, for different $N$ and service packet periods. Jitter refers to the variation in packet transmission latency over time within a network. It is the inconsistency observed in the arrival time of packets, primarily caused by varying network congestion levels, route changes, or other unforeseen factors. To obtain the median and average results, we generated $100$ scenarios for each packet order point. The median offers insight into the typical experience of a node without the influence of outliers, while the average jitter aggregates these variations to provide a general sense of the network’s jitter over time. According to Fig. 15, MSF shows an initial peak (median and average jitters), due to the network’s formation phase, which might disrupt real-time services. In contrast, PB consistently maintains lower jitter performances than MSF, which indicates a stable packet delivery time. The median jitter of MSF exhibits high variability, thus unsuitable for services requiring consistent data rates, while PB follows a more predictable jitter performance, which makes it fit for real-time applications. As $N$ increases from $50$ to $100$, and for a given scheme and service packet period, we notice that the jitter performance is stable and does not degrade. For instance, for a service packet period of $5$ seconds and PB scheme, the average jitter is at 0.6 seconds for $N=50$ and $N=100$. However, we notice that for $N=50$, variations in median jitter is slightly higher than for $N=100$. This might be due to the fixed slotframe length that might not be suitable for any network size. For a given $N$, the jitter degrades with the service packet period. Indeed, a longer service time impacts packet transmissions and thus might cause additional delay and/or instability. In summary, the jitter is slightly impacted by the size of the network, while the influence of the service packet delay is more significant on both the average and median jitters.
@@
Fig. 16 presents the Cumulative Distribution Function (CDF) as a function of the charge consumed (in Coulomb) for the PB (blue) and MSF (purple) solutions, for different $N$ and service packet periods. For any scenario, the charge consumption profiles of PB and MSF are almost identical. This result emphasizes the efficiency of our solution in conserving energy similarly to MSF, while realizing significant improvements in terms of other performance metrics, such as joining time, E2E latency, and jitter. We note that for a given $N$, when the service packet period increases, the CDF’s slope is almost the same. This means that the amount of charge consumed by a node in the network is not affected by the service packet period. Also, the CDF between $N=50$ ad $N=100$ at a fixed service packet period are very similar. Hence, $N$ has a negligible impact on the consumed charge per node. In Fig. 17, we evaluate the CDF of the lifetime of the network’s nodes for the PB (blue) and MSF (purple) solutions, for different $N$ and service packet periods. For any $N$ and service packet period value, the lifetime of nodes in the PB-based network is slightly lower than that of the MSF-based one. For instance, for $N=50$ and service packet period equal to $45$, we see that $67\%$ of the nodes will have a lifetime of less than 5 years using PB, while this proportion is only $58\%$ for MSF. When $N$ increases, the gap between PB and MSF gets lower, which highlights the capability of our method to save energy in dense networks.
V Conclusion
Our research introduces a novel approach to enhance 6TiSCH network performances by carefully optimizing the interaction between SF and RPL. Through a multitude of modifications within the exchanged messages in 6TiSCH and improvements in SF and RPL operations, our proposed solution is capable of enhancing several performance metrics, including joining time, transmitted/received packet number, E2E latency, and jitter, while keeping the energy usage at levels similar to the MSF benchmark and nodes’ lifetime. Despite the realized improvements, several challenges remain to be solved. For instance, the fixed duration of time slots prevents the network from achieving its full potential.

### v2

@@ 4.4 Scheduling Packet Update
In the context of TSCH, transmitting packets from higher-layer protocols is not an instantaneous process; rather, it entails a meticulous queuing and scheduling mechanism. As depicted in Fig. 9, incoming packets are first enqueued within the TSCH queue and await their transmissions. Since we inject the list of time slots in DIO and DAO packets, they will be locked during time $T_{q}=T_{\rm de}-T_{\rm en}$, where $T_{\rm en}$ is the instant at which the packet entered the TSCH queue, and $T_{\rm de}$ is the time at which it left the queue. To avoid locking the list of time slots for a long time $T_{q}$, we propose to inject them within the packet only before leaving the queue, i.e., at instant $T_{de}$, hence $T_{q}\approx 0$. By carefully orchestrating the RPL control packets’ update, through adding details about the schedule state (i.e., the free and reserved slot IDs), we optimize time slot utilization.
@@
4.5 Avoiding TSCH Queue Overflow
@@
4.6 Adaptive Self-Sufficient Cell Reservation
In a typical 6TiSCH network, all DAO packets are destined towards the DAG root, thus navigating in a multi-hop fashion. In our proposal, essential information about cell reservation, which consists of the available slots and proposed slot lists, is embedded within these DAO packets. When a parent node receives a DAO packet, as it is the principal conduit to the DAG root, it reserves new cells for its new child where the number of cells should not exceed the maximum number of cells to allocate at once. Then, it loads the list of its free slots (instead of the old one that belongs to the child) and the number of slots to be selected (slots to be allocated to its new child) into the DAO packet. This packet is then seamlessly forwarded, with each intermediate node towards the DAG root executing the same procedure. As shown in Fig. 10, upon reception of a DAO packet from node $3$, node $i$ undertakes cell selection for reservation and embeds the list of chosen slots into the TSCH ACK packet. It then forwards the enhanced packet to its parent.
@@ 4.6 Adaptive Self-Sufficient Cell Reservation
To reduce latency at every hop, we propose that the receiver of a DAO packet selects the nearest time slot to the actual one used for the reception. One way to do it is to rely on the concept which is similar to the concept of the Low-Latency SF (LLSF) protocol [32], where the slots are selected according to the gap to the left. However, we propose here selecting the nearest slot to the actual one in any direction. Accordingly, if the node is in the process of joining the DODAG, it selects the nearest free slot to the actual one to start its activity with minimal delay. This mechanism is equally effective for nodes that have already joined, as they continue using the slot allocated to their parent, ensuring better slot continuity, reducing scheduling overhead, and improving end-to-end transmission efficiency. Consequently, both energy and latency tied to time slot reservations can be reduced.
5 Performance evaluation
5.1 Simulation Configuration
To comprehensively assess the efficiency and resilience of our proposed scheme, we employed the 6TiSCH simulator [33]. This simulator provides a realistic environment for analyzing key network performance metrics, including packet transmission success rates, queue behavior, congestion management, energy consumption, and parent switching efficiency [7]. Also, this tool has been extensively used in previous studies to validate 6TiSCH-based scheduling and routing mechanisms [7, 34, 35, 36]. For the sake of thorough analysis, generated plots include outlier results, thus reflecting the full range of observed results.
Our simulations spanned $30$ minutes and $270$ minutes ($4.5$ hours) and were repeated 100 times across two distinct network configurations:
@@ 5.1 Simulation Configuration
We conduct our evaluation by randomly placing $N_{1}=50$ (resp. $N_{2}=100$) nodes within a 1 km2 area for each simulation run. The $x$ and $y$ coordinates of these nodes are generated randomly. We employed the Pister-hack connectivity model [33], where each node maintains at least three links to neighbors. Throughout the simulations, the observed routing trees ranged from a depth of $2$ hops to $3$ hops, with a median depth of $2$ hops for the network of $N_{1}=50$ nodes (resp. a median depth between $2$ and $3$ for the network with $N_{2}=100$ nodes). Moreover, we assume that the frequency of data transmissions for the application or service is in the values $\{5,15,45,60\}$ seconds.
The remaining simulation parameters are presented in Table 1.
Table 1: Simulation parameters
Parameters Value
Number of nodes 50, 100
Number of simulation runs 100
Simulation duration (min) 30, 270
Network area (km2) 1
Service or application packet period (sec) 5, 15, 45, 60
Slotframe length (in number of slots) 100
Time slot duration (sec) 0.01
Secure Joining required True
TSCH queue size (number of packets) 10
Number of radio channels 16
TSCH maximum payload length (bytes)
DIO packet size (bytes)
DAO packet size (bytes)
MSF PB
90 120
76 120
20 75
5.2 Parameters’ Values Setup
5.2.1 DIO and DAO Packet Size Selection
Based on eq.(4), with a slotframe size $Slotframe_{Length}=100$ slots, the largest size of the free slots’ list is 47 slots. Since each slot ID is encoded on 1 byte, this requires adding 47 bytes to a packet. To accommodate this additional information while ensuring the packets remain within the TSCH maximum payload length (i.e., 120 bytes), we extend the conventional DIO and DAO packet sizes from 76 and 20 bytes to 120111With the additional 44 bytes for the DIO, a maximum of 44 slot IDs can be injected. and 75 bytes, respectively.
5.2.2 Queue Management Parameters
Upon each packet enqueuing operation, nodes assess the remaining capacity within the TSCH queue. When the available space in the queue is less than or equal to “TSCH queue empty places threshold”, the node initiates a cell reservation request using the 6P protocol. It will request “Number of cells reserved upon overflow” cells, and subsequent request is constrained by the value of the parameter “Cycle interval before next cell request”.
A low value of “TSCH queue empty places threshold” may prevent the mechanism from functioning effectively. Conversely, a higher threshold may lead to extensive cell reservation requests, thus increasing the number of allocated cells beyond what is necessary. While this may accelerate packet transmission, it can also result in inefficient resource utilization. The delay of “Cycle interval before next cell request” is essential to avoid multiple consecutive reservation requests. However, its value must be carefully balanced. A low value may not allow enough time to resolve the previous overflow before triggering another request, leading to multiple cell reservation requests, which can increase scheduling overhead and inefficient resource allocation. In contrast, a long period may cause the node to overlook queue overflows, delaying necessary cell reservations and potentially increasing packet loss.
@@ 5.2.2 Queue Management Parameters
Figure 11: Number of dropped packets due to queue overflow vs. size of “Cycle interval before the next cell request”.
To define the value of parameter “Cycle interval before next cell request” and demonstrate the performance of our mechanism, we conducted simulations with an application period of 5 seconds. This setup models high-traffic conditions, where queue congestion is most significant and the effectiveness of queue management mechanisms becomes more evident. In Fig. 11, the performance of the MSF benchmark is compared with our proposed solution under two configurations, namely $PB\_2$ and $PB\_5$, representing cycle intervals of 2 and 5 slotframes, respectively. As shown, MSF exhibits a significantly higher number of packet drops due to queue overflow, with notable variability. In contrast, our method, using $PB\_2$ or $PB\_5$, substantially reduces packet losses. With $PB\_2$, our technique demonstrates slightly better performances than for $PB\_5$, highlighting that a cycle interval of 2 slotframes is sufficient to mitigate queue overflow. Increasing the interval beyond 2 slotframes would delay necessary slot allocations, potentially impacting data transmission efficiency. Hence, in the remaining, we select Cycle interval before next cell request” to be equal to 2 slotframes.
(a) Joining time (in sec)
(b) Consumed current (in $\mu$A)
Figure 12: Joining time and consumed current vs. nbr. of proposed slots per DIO.
5.2.3 Number of Proposed Slots per DIO
To set the parameter “Number of proposed slots per DIO” to the best value, we conducted several experiments with different values. Specifically, in Figs. 12(a)-12(b), we evaluated the impact of different numbers of proposed time slots per DIO on the joining time and current consumption (equivalent to energy consumption).
These experiments have been conducted over an observation time of $30$ minutes, with an application period of $5$ seconds, and for different network sizes, i.e., $N_{1}=50$ and $N_{2}=100$ nodes. As shown in Fig. 12(a), for any number of nodes, the joining time reduces when the number of proposed slots is higher. This is expected since more proposed slots mean that a novel node can rapidly be configured and join the DODAG. However, since the number of proposed slots cannot be very high, as it may increase the size of packets, our preference is for the value $7$ since it provides similar results to the system with the value $13$, in particular for $N_{2}$, i.e., for a large network. A similar explanation can be provided for Fig. 12(b) where the value $7$ achieves low current consumption for the large network. Indeed, a very high value, such as $13$, can result in inefficient idle listening, thus increasing power consumption. Consequently, we recommend the “number of proposed slots per DIO” to be $7$ as the best value since it is the lowest one that achieves good performances in terms of joining time and consumed power.
(a) Joining time (in sec)
(b) Consumed current (in $\mu$A)
Figure 13: Joining time and consumed current vs. proposed DIO cells duration.
5.2.4 DIO Cell Duration Selection
The selection of the “DIO cells duration” parameter (expressed in number of slotframes), which defines how long a node maintains proposed cells for new child nodes in its DIO, is based on extensive simulation experiments following the same methodology used in Subsection 5.2.3. The DIO cells duration was tested for different values, i.e., 3, 7, 10, 13, and 17 slotframes, while measuring key performance metrics, i.e., join time and current consumption, in different network sizes $N_{1}=50$ and $N_{2}=100$ nodes, and illustrated in Figs. 13(a)-13(b). These metrics were assessed as this mechanism aims to reduce the joining time while minimizing energy drain. According to Fig. 13(a), the joining times are similar, e.g., around 10.5 seconds for $N_{1}$, for any size of DIO cells duration. However, when looking at Fig. 13(b), we notice that energy consumption is the smallest for a duration equal to 10 slotframes. This suggests that a DIO cells duration of 10 slotframes provides the best trade-off between reducing joining time and minimizing energy consumption.
We summarize in Table 2 the selected parameters for our solution.
Table 2: Parameters of the proposed solution
Parameter Value
Number of proposed slots per DIO 7
DIO cells duration (in number of slotframes) 10
Number of permanent DIO slots 1
Slot selection ratio within DIO slots 3
Maximum number of cells to reserve at once 5
Initial phase duration (min) 45
Queue Management Parameters
TSCH queue empty places threshold 2
Cycle interval before next cell request 2
Number of cells reserved upon overflow 1
@@ 5.2.4 DIO Cell Duration Selection
Figure 14: Joining time (in sec) vs. length of service packet period (short term).
5.3 Simulation Results
For our simulations, we define two scenarios, the first is in the short term with a duration of $30$ min, while the second is in the long term with a duration of $4.5$ hours.
5.3.1 Short Term Results
The initial phase of the network’s life is known as network formation. This phase is pivotal for establishing the network’s parameters and topology.
To demonstrate the effectiveness of our solution in accelerating the network formation timeline and minimizing power consumption, we conducted simulations that last $30$ minutes.
@@ 5.3.1 Short Term Results
Figure 15: Consumed current (in microampere) vs. length of service packet period (in sec) (short term).
@@
Figure 16: Nbr. of Tx/Rx packets (different service packet periods, long term).
As shown in Fig 14, our proposed solution PB achieves lower joining time for different number of nodes within the network, $N$, and service packet period duration. For instance, PB achieves an average joining time of $11$ seconds for a service packet period equal to $15$ and $N=50$, while MSF needs approximately $12$ seconds in the same conditions. Indeed, a node within the PB framework uses proposed slots in the received DIO to define the cells during which it sends the DAO, unlike MSF which relies on sending the DAO packet and the 6P request to its preferred parent and then waiting for the 6P response for cell reservation. In PB, the DAO packet serves a dual purpose: Communicating with the RPL to establish the DODAG structure and participating in SF. As the service packet period increases, we notice a degradation in the joining time. This is expected since a node with a low (resp. high) service packet period will transmit EB packets more (resp. less) often, which increases (resp. reduces) the likelihood that other nodes synchronize and join the DODAG.
In 6TiSCH networks, the process of network formation often expends a significant amount of energy, even before useful data transmission. To evaluate energy consumption, we adopt a model based on the operational state of nodes in each time slot, as implemented in the 6TiSCH simulator. Time slots in which the node is desynchronized are not considered as the node is inactive in such cases. The total charge consumption over the simulation period $Q_{\text{total}}$ is computed as the sum of the charge consumed in each state $C_{\text{state}}$ [37], i.e.,
$Q_{\text{total}}=\sum_{\text{state}}N_{\text{state}}\cdot C_{\text{state}},$ (5)
where $N_{\text{state}}$ represents the number of occurrences or total duration spent in each state. In Fig. 15, we present the consumed current related to the scenarios of Fig. 14. The PB scheme performs better than the MSF one, for any $N$ and service packet period duration. The performance gap increases as $N$ becomes higher. The observed improvement in join time and energy efficiency are due to embedding slot proposals within the DIO packet rather than using 6P transactions. Also, by merging control packets, our method accelerates node activation, enabling newly joined nodes to quickly invite other nodes into the DODAG and enhancing the overall network convergence. As the network scales up, competition for shared resources intensifies leading to increased transmission attempts and current consumption. However, our solution remains particularly robust in such conditions. For instance, in Fig. 15, MSF exhibits high current consumption compared to PB in any network size. The ability of PB to proactively allocate resources results in more stable current consumption. Finally, we notice that the service packet period has a slight impact on the total consumed current.
5.3.2 Long term results
We assess here the performances of PB and MSF for long-term operation, in terms of the number of transmitted/received packets, end-to-end (E2E) latency, network jitter, consumed charge, and nodes’ lifetime.
Fig. 16 illustrates the number of transmitted (Tx) and received (Rx) packets for different service packet periods and $N$. For any $N$ and service packet period, our solution outperforms the MSF benchmark. Indeed, the network that uses our approach realizes more frequent transmissions, and nodes within it successfully receive more packets. When the service packet period increases, i.e., a longer service is running, the number of exchanged packets drops. such a result is expected since during the observation time of 4.5 hours, a longer service time impacts the frequency of its execution over the observation period. Moreover, when $N=100$, the number of exchanged packets almost doubles compared to the case with $N=50$. Indeed, the availability of more nodes joining the DODAG makes it possible to transmit/receive more packets, within the limits of available channels and time slots of 6TiSCH.
@@ 5.3.2 Long term results
Figure 17: E2E latency (in sec) vs. service packet period (long term).
End-to-end latency is defined as the duration measured from the absolute slot number (ASN) at which a source node transmits a service/application packet, to the ASN at which this packet is successfully received and acknowledged by the destination, while jitter refers to the variation in packet transmission latency over time within a network. It is the inconsistency observed in the arrival time of packets, primarily caused by varying network congestion levels, route changes, or other unforeseen factors. Fig. 17 shows the network’s E2E latency for different service packet periods and $N$. As it can be seen, the PB scheme achieves lower E2E latency in any network condition than the MSF. This result underscores the efficiency of our method.
(a) $N=50$
(b) $N=100$
Figure 18: Median and avg. jitter of Tx packets (different service packet periods, long term).
In Fig. 18, we plot the median and average jitter performance for the transmitted packets during the observation time, for different $N$ and service packet periods. To obtain the median and average results, we generated $100$ scenarios for each packet order point. The median offers insight into the typical experience of a node without the influence of outliers, while the average jitter aggregates these variations to provide a general sense of the network’s jitter over time. According to Fig. 18, MSF shows an initial peak (median and average jitters), due to the network’s formation phase, which might disrupt real-time services. In contrast, PB consistently maintains lower jitter performances than MSF, which indicates a stable packet delivery time. The median jitter of MSF exhibits high variability, thus unsuitable for services requiring consistent data rates, while PB follows a more predictable jitter performance, which makes it fit for real-time applications. As $N$ increases from $50$ to $100$, and for a given scheme and service packet period, we notice that the jitter performance is stable and does not degrade. For a given $N$, the jitter degrades with the service packet period. Indeed, a longer service time impacts packet transmissions and thus might cause additional delay and/or instability. In summary, the jitter is slightly impacted by the size of the network, while the influence of the service packet delay is more significant on both the average and median jitters.
@@
Figs. 19 and 20 present the Cumulative Distribution Function (CDF) of charge consumption and node lifetime, respectively, for PB and MSF under different network sizes ($N=50$, $N=100$) and service packet periods (5, 15, 45, 60s). The node lifetime is determined by
Lifetime=Battery Capacity (µAh)Average Current Consumption ($\text{Lifetime}=\frac{\text{Battery Capacity (µAh)}}{\text{Average Current Consumption (µA)}\times 24\times 365},$ (6)
where the “Average current consumption” corresponds to the mean current consumption trends presented in Fig. 19. The CDF in Fig. 19 shows that PB and MSF consume nearly identical charge per node across all scenarios, confirming that PB does not introduce additional energy overhead. However, the CDF in Fig. 20 reveals that PB nodes exhibit slightly shorter lifetimes than MSF nodes, despite having the same total charge consumption. This suggests that PB nodes experience a higher current consumption per time unit due to increased activity, implying that PB nodes spend less time in low-power states, such as idle listening, sleep, or desynchronization. For $N=100$, the difference in node lifetime between PB and MSF becomes less pronounced, indicating that PB uniformly distributes energy across nodes in dense networks. The similarity in charge consumption across different network sizes confirms that PB maintains stable energy consumption.
Simulation results demonstrate that the proposed solution consistently outperforms the MSF benchmark in terms of end-to-end latency, jitter, number of transmitted/received packets, and energy efficiency across both short-term and long-term scenarios. Notably, our solution maintains stable performance as the network size increases from $N=50$ to $N=100$ nodes, highlighting its scalability for large IIoT networks. The cumulative energy consumption remains consistent with increasing node density, indicating preserved efficiency under varying network conditions. These findings prove our method’s suitability for real-time and energy-constrained applications and its effectiveness in large and dynamic IIoT systems.
6 Discussion and Future Directions
Despite the demonstrated advantages, the proposed approach presents certain trade-offs and limitations. For instance, our slot merging strategy is limited by the maximum payload size imposed by TSCH, which must be respected to avoid packet fragmentation. Exceeding this limit leads to fragmentation, thus increasing transmission time, energy consumption, and control overhead while undermining the benefits of packet merging by creating additional fragments that must be transmitted separately. Consequently, fragmentation can cause delays, raise the risk of packet loss, and negate the intended efficiency gains. Moreover, when the size of the injected data grows, the processing time required for slot selection increases. To ensure that slot selection is completed within the TSCH slot duration, it may be necessary to adjust the latter’s duration according to the device’s computational capability. However, increasing the slot duration reduces the overall network throughput, hence creating a trade-off between scheduling efficiency and transmission capacity. This challenge underscores the need for optimized slot selection algorithms that minimize processing time without extending the TSCH slot duration.
Also, in the proposed approach to manage TSCH queue overflow, knowing the reason behind packet accumulation enables more informed decisions, greater efficiency, and increased resilience against edge-case scenarios. Indeed, initiating a reservation request without identifying the underlying cause could worsen the situation. For example, if packet accumulation results from a temporary disconnection with the parent node (disjoin), sending a reservation request would be inefficient and may lead to unnecessary resource consumption. Conversely, if the accumulation is due to a lack of available transmission slots or network congestion, requesting additional cells becomes suitable. Therefore, in our future work, we will focus on developing a decision-making mechanism that diagnoses the cause of packet accumulation to determine the appropriate mitigation strategy.
This work serves as a prototype implementation of the proposed approaches, enabling the optimization of key parameters and analysis of their impact on network efficiency. However, transitioning from simulations to real-world deployments introduces additional challenges, such as hardware limitations, environmental interference and obstacles, multi-protocol co-existence, and scalability, which warrant further investigation. In future work, we aim to validate the proposed approach through hardware-based experiments and explore alternative slot encoding methods to tackle time slot duration and TSCH payload constraints, which would enhance the proposed solution’s scalability and adaptability in real environments.
Finally, there is a correlation between the slotframe length, TSCH payload constraints, and time slot duration. This relationship depends on several factors, including the network size and the node’s physical characteristics. In future work, we will investigate this relationship and develop strategies for optimal scheduling, data compression, and adaptive slotframe length configuration to maximize scheduling efficiency.
7 Conclusion
In this paper, we introduced a novel approach that optimizes the performances of the 6TiSCH network. Specifically, through the integration of several proposed changes within the exchanged control messages and operations of SF and RPL, our solution improves network formation by reducing node joining time and energy consumption. Moreover, it reduces end-to-end latency and jitter, while enabling a higher number of successfully received packets compared to the MSF benchmark. Also, it keeps energy usage at levels similar to those of the MSF benchmark. This work provides important insights into the enhancement of 6TiSCH operations for future 5G and beyond IIoT networks. Nevertheless, challenges remain to unleash its full potential, including balancing slotframe length and overcoming the TSCH payload limit. Future work will focus on adaptive scheduling for dynamic slotframe adjustment, intelligent queue management to avoid unnecessary cell reservations, and real-world experimental validation to assess performances under practical conditions.